# Inertial Coordination Games

### v2

@@ 3   Equilibrium law of motion
Hence, the limit belief threshold $\mu_{\infty}^{*}$ pins down limit aggregate play at state $\theta$. In particular, $\lambda_{\infty}(\theta)=RD(\theta)$ a.e. if and only if $\mu_{\infty}^{*}=c-(1/2)$. In words, the risky action is adopted by the whole population in the limit if and only if, in the complete information game in which the fundamental $\theta$ is known, playing the risky action is the best response to the uniform conjecture that $1/2$-proportion of players are playing the risky action. Hence, we call $c-(1/2)$ the risk dominance threshold.
@@
which implies $\lambda_{t}=1$. The argument for $\theta+\lambda_{0}<c$ is analogous so $\lambda_{t}=0$ for all $t\in\mathcal{T}$. Hence, limit play depends on $\lambda_{0}$: $\lambda_{\infty}(\theta|\Sigma,\lambda_{0})=\mathbbm{1}[\theta\geq c-\lambda_{0}].$$\diamondsuit$
@@
What are we to make of these differing predictions? In what follows we offer a unified account of when risk-dominance or history-dependence obtains as a function of the learning process. It will turn out that Example 1 is a special case of slow learning since posterior precisions grow sub-quadratically (in Example 1, they do not change beyond $t=1$). On the other hand, Example 2 is a special case of fast learning since posterior precisions are super-quadratic (in Example 2, they “jump immediately to infinity” at time $t=1$). [^27] Translated into a static incomplete information game, this correspond to interspersing interim Bayesian updating with rounds of interim deletion of strictly dominated strategies. Online Appendix V makes this connection precise.
@@ 4   Characterization of limit play
[…]
@@
Putting parts (i) and (ii) of Theorem 1 together offers simple but novel economic implications for how the ex-ante joint distribution over shocks $(\lambda_{0})$ and fundamentals ($\theta$) interact with learning speeds to shape limit outcomes. Figure 3b depicts the regions of realizations of shocks and fundamentals under which shocks propagate so the limit play is the risky action (red region), or fizzle out so limit play is the safe action (blue region). Panel (a) depicts this for the case of sub-quadratic learning. As Theorem 1 (i) establishes, the risk-dominant boundary is selected so limit play is history independent, and does not depend on the shock’s realization. Panel (b) depicts the case of super-quadratic learning which rotates the boundary clockwise around the point $(1/2,c-1/2)$. [^30] We have depicted this so that the state selects initial-play dominance i.e., $\lambda_{\infty}=1\iff\theta\geq c-\lambda_{0}$. We will later establish (Proposition 2 that initial-play dominance and risk-dominance span possible limit outcomes implementable via any learning speed $(\sigma_{t})_{t}$. More generally, for a given super-quadratic learning rate, this will lead to a (potentially non-linear) boundary between these two extremes. Hence, when learning is fast, the regime might now survive when the state is weak as long as the initial shock is small (region $A$); on the other hand, the regime might now also fail when the state is strong when the initial shock is large (region $B$).
@@
Figure 4 illustrates the forces underlying Theorem 1 (i) for the case where $\lambda_{0}<1/2$. Each panel corresponds to a different time. In each panel, the dotted line represents the state $\theta$, and the red curve is the distribution of posterior means $\mu_{t}$. The blue line is the threshold belief $\mu_{t}^{*}$, and the shaded area is the mass of agents $\lambda_{t}$ who take the risky action because their beliefs exceed the threshold. As time passes, agents learn and $\mu_{t}$ contracts around $\theta$. In Figure 4, learning and hence the speed of this contraction is slow. Consequently, the evolution of $\mu_{t}^{*}$ is relatively fast. As $t\to\infty$, $\mu_{t}^{*}$ converges to $c-(1/2)$ and $\mu_{t}$ converges to a point mass on $\theta$ due to limit learning. Since $\theta\geq c-(1/2)$, aggregate play converges to the risk dominant (risky) action.
Figure 5 illustrates the forces underlying Theorem 1 (ii) for the case where $\lambda_{0}<1/2$. With super-quadratic learning, the speed of contraction of $\mu_{t}$ is fast. Consequently, the evolution of $\mu_{t}^{*}$ is relatively slow. As $t\to\infty$, $\mu_{t}^{*}$ converges to a limit threshold $\mu_{\infty}^{*}$ that is bounded away from $c-(1/2)$, and since $\theta\in(c-(1/2),\mu_{\infty}^{*})$, aggregate play converges to the non-risk dominant (safe) action.
@@
In particular, for any threshold $\mu^{*}$ between $c-\lambda_{0}$ and $c-(1/2)$, there exists some learning process $\Sigma\geq 0$ such that:
[…]
@@ 5   Transition dynamics
We have thus far focused on limit play. We now turn to studying the connection between the speed of learning and the time path of aggregate play. We will consider the iid signal case such that the learning process $\Sigma(\sigma)$ is parametrized by the standard deviation of each period’s signal $\sigma>0$. We focus on the iid case because it is a canonical model of learning, and makes results easier to state since $\Sigma$ is now parametrized by a scalar; nonetheless, the qualitative features of our results extend beyond this case. Within this iid learning environment, recall that Corollary 2 (i) implies that the limit action is risk dominant for all fundamentals and initial conditions. The following result shows that the path of aggregate play exhibits qualitatively different behavior depending on learning rates.
@@
Fix any $\lambda_{0}\in(0,1)$ and $\theta$ between $c-1/2$ and $c-\lambda_{0}$. Define $\overline{\beta}:=1+|c-(1/2)-\theta|/2$. For any $\epsilon>0$, $\alpha\in(0,1)$, and $\beta\in(1,\overline{\beta})$, there exists $\underline{\sigma},\overline{\sigma}>0$ such that:
@@
(b) $\sigma=0.01$: sudden transition
Proposition 3 is subtly connected to the literature on coordination games. A common theme in this literature is that heterogeneity—in preferences, beliefs, etc.—across players can induce uniqueness in static supermodular games (Morris and Shin, 2006). A dynamic version of this insight is that with lots of hetrogeneity, as aggregate incentives shift, the set of equilibria vary smoothly—this is analogous to Part (i) of Proposition 3; by contrast, with little heterogeneity, the set of equilbiria can vary discontinuously—this is analogous to Part (ii) of Proposition 3. The key insight of Proposition 3 is that learning rates induce time-varying hetrogeneity in beliefs which, in turn, matters for transition dynamics. [^33] Koh and Morris (2022) relates heterogeneity to the speed of contagion in a networked population. This can be equivalently interpreted (see, e.g., Morris (1997)) as best-response dynamics or interim deletion of strictly dominated strategies where each player’s beliefs and higher-order beliefs are encoded by the network. However, Koh and Morris (2022) study speed of contagion for graphs/beliefs which are held fixed i.e., in the incomplete information game interpretation, there is no learning.
@@ 6   Concluding Remarks
We have developed a simple but novel model of inertial coordination games. This allowed us to combine and unify insights from the literature on best-response dynamics in complete-information coordination games (Crawford, 1995) and global games (Carlsson and Van Damme, 1993). We think that inertial coordination games serve as a natural description of many ‘real world’ coordination environments with both shrinking uncertainty about fundamentals (e.g., solvency of a bank, strength of a currency regime, quality of a networked product) as well as an endogenously changing state shaped by past play (e.g., past withdrawals, short positions, past sales). At the same time, we also showed that inertial coordination games share much of the game-theoretic logic underpinning coordination and learning in games. [^34] We show in Appendix B that for a fixed speed of learning (as a function of physical time), limit equilibrium predictions coincide as inertia vanishes (gap between time periods shrink). In this regard, inertial coordination games generalizes the setting with contemporaneous coordination motives. Nonetheless for any fixed level of inertia, the sub/super-quadratic boundary continues to hold. In this regard, we offer a different perspective on ‘history versus expectations’ (Krugman, 1991) in economic modeling—in the presence of inertia, expectations about history shape coordination outcomes. In particular, the speed at which players grow confident is a crucial determinant for whether coordination succeeds—the risk-dominant action is played in the limit if and only if posterior precisions grow sub-quadratically (Theorem 1). This delivered a suite of rich predictions of whether and to what extent initial shocks can propagate. In the super-quadratic regime, we showed that any profile of limit play between risk- and initial play-dominance can be rationalized by some learning speed. In the sub-quadratic regime, we further identified qualitatively different behavior in the path of aggregate play (Proposition 3): when learning is slow, transition dynamics are gradual; when learning is fast, sudden.
@@ A.1  System Dynamics
Conditional on $\theta$, the time-$(t-1)$ distribution of posterior means is $\mu_{j,t-1}|\theta\sim N(\theta,\eta_{t-1}^{2})$. Since agents’ information sets are independent conditional on $\theta$, this implies that, under agent $i$’s time-$t$ posterior belief, the marginal distribution of time-$(t-1)$ posterior means is $\mu_{j,t-1}|\mathcal{F}_{it}\sim N(\mu_{it},\eta_{t}^{2}+\eta_{t-1}^{2})$. Hence, by the inductive step, $i$ takes $a_{it}=1$ if and only if
@@
Let $F(\mu_{it})$ be the RHS function. Since $F$ is continuous and strictly increasing in $\mu_{it}$, and $F(c)>c$ and $F(c-1)<c$ hold, $F(\mu_{it})=c$ has a unique solution $\mu_{t}^{*}\in(c-1,c)$. Hence, there exists a unique cutoff $\mu_{t}^{*}$, defined implicitly by ($\mu^{*}$-LoM) at time $t$, such that $i$ takes $a_{it}=1$ if and only if $\mu_{it}\geq\mu_{t}^{*}$.
Finally, since $\mu_{it}|\theta\sim N(\theta,\eta_{t}^{2})$, we see that
$\lambda_{t}(\theta)=\mathbb{P}_{\theta}(\mu_{it}\geq\mu_{t}^{*})=\Phi\left(\frac{\theta-\mu_{t}^{*}}{\eta_{t}}\right)$
and hence ($\lambda$-LoM) holds at time $t$, as desired.
@@
Assume $\mu_{t-1}^{*}=c-(1/2)$. Since $\mu_{t}^{*}:=c-(1/2)$ solves ($\mu^{*}$-LoM) and it has a unique solution, $c-(1/2)$ is a steady state. To see that it’s the unique steady state, assume $\mu_{t-1}^{*}=\mu_{t}^{*}$. Then, ($\mu^{*}$-LoM) reduces to $\mu_{t}^{*}+\Phi(0)=c$ and hence $\mu_{t}^{*}=c-(1/2)$, as desired.
@@
$\mu_{t+1}^{*}>\mu_{t}^{*}$
[…]
@@
    $|\mu^{*}_{t+1}-\mu_{t}^{*}|$ is decreasing in $t$ [^35] Note that parts (i),(ii), and (iii) of this Lemma hold without Assumption 1.
@@
Part (iii). Note that ($\mu^{*}$-LoM) implies
$|\mu^{*}_{t+1}-\mu_{t}^{*}|=\sqrt{\eta_{t}^{2}+\eta_{t+1}^{2}}\left|\Phi^{-1}(c-\mu_{t+1}^{*})\right|.$
Note that $\sqrt{\eta_{t}^{2}+\eta_{t+1}^{2}}$ is non-increasing in $t$. First, assume $\lambda_{0}>\frac{1}{2}$. Then, $(c-\mu_{t+1}^{*})>\frac{1}{2}$ decreasing in $t$ implies $\Phi^{-1}(c-\mu_{t+1}^{*})>0$ is decreasing in $t$, and therefore $|\Phi^{-1}(c-\mu_{t+1}^{*})|$ is decreasing in $t$. Hence, $|\mu^{*}_{t+1}-\mu_{t}^{*}|$ is also decreasing in $t$.
@@
Part (iv). Wlog, assume $\lambda_{0}>\frac{1}{2}$. First, we show $(\theta-\mu_{t}^{*})/\eta_{t}\to-\infty$ when $\theta<\mu_{\infty}^{*}$. Since $((\theta-\mu_{t}^{*})/\eta_{t})_{t}$ is decreasing and becomes negative for sufficiently large $t$, it should either converge to some negative constant or diverge to $\infty$- ∞. Since $\lim_{t\to\infty}(\theta-\mu_{t}^{*})<0$, $\lim_{t\to\infty}(\theta-\mu_{t}^{*})/\eta_{t}=-\infty$ must hold by limit learning.
@@
Part (i). Subtracting ($\mu^{*}$-LoM) at time $t-1$ from ($\mu^{*}$-LoM) at time $t$ yields:
@@
Wlog, assume $\lambda_{0}>(1/2)$, then $\gamma_{1}>0$. Suppose for a contradiction that $\gamma_{2}<0$: then $\gamma_{1}<\gamma_{2}<0$, which is a contradiction since $\gamma_{1}>0$ by definition. Hence, $\gamma_{1}\geq\gamma_{2}\geq 0$. By induction, $(\gamma_{t})_{t}$ is a monotone decreasing sequence bounded below by 00, so it converges: $\gamma_{t}\downarrow\gamma_{\infty}\geq 0$, as desired. ∎
@@
Part (i). Fix any $x>0$. For any $A^{\prime}>A>0$, note that
$A^{\prime}F_{A^{\prime}}(x)-AF_{A}(x)=\Phi(F_{A}(x))-\Phi(F_{A^{\prime}}(x))$
Suppose for a contradiction that $F_{A^{\prime}}(x)\geq F_{A}(x)$. Since $\Phi$ is increasing, the above implies
$A^{\prime}F_{A^{\prime}}(x)\leq AF_{A}(x)$
which is a contradiction since $A^{\prime}>A$.
@@
$Ay^{\prime}=\varphi(x)-\varphi(y)y^{\prime}$
$\implies y^{\prime}=\frac{\varphi(x)}{A+\varphi(y)}>0,$
@@
Then $A_{\gamma_{1}}(x)\uparrow\bar{A}^{*}$ follows from continuity of $A_{\gamma_{1}}(\cdot)$.

### v3

@@ 3 Equilibrium law of motion
Hence, the limit belief threshold $\mu_{\infty}^{*}$ pins down limit aggregate play at state $\theta$. In particular, $\lambda_{\infty}(\theta)=RD(\theta)$ a.e. if and only if $\mu_{\infty}^{*}=c-(1/2)$. In words, the risky action is adopted by the whole population in the limit if and only if, in the complete information game in which the fundamental $\theta$ is known, playing the risky action is the best response to the uniform conjecture that $1/2$-proportion of players are playing the risky action. Hence, we call c(1/2)c-(1/2) the risk dominance threshold.
@@
[…]
@@
What are we to make of these differing predictions? In what follows we offer a unified account of when risk-dominance or history-dependence obtains as a function of the learning process. It will turn out that Example 1 is a special case of slow learning since posterior precisions grow sub-quadratically (in Example 1, they do not change beyond $t=1$). On the other hand, Example 2 is a special case of fast learning since posterior precisions are super-quadratic (in Example 2, they “jump immediately to infinity” at time t=1t=1). [^27] Translated into a static incomplete information game, this correspond to interspersing interim Bayesian updating with rounds of interim deletion of strictly dominated strategies. Appendix V makes this connection precise.
@@ 4 Characterization of limit play
    $\eta^{-2}_{t}=O(t^{2-\epsilon})\implies\lambda_{\infty}(\theta)=RD(\theta)\quad\text{a.e.}$
@@
Putting parts (i) and (ii) of Theorem 1 together offers simple but novel economic implications for how the ex-ante joint distribution over shocks $(\lambda_{0})$ and fundamentals ($\theta$) interact with learning speeds to shape limit outcomes. Figure 3b depicts the regions of realizations of shocks and fundamentals under which shocks propagate so the limit play is the risky action (red region), or fizzle out so limit play is the safe action (blue region). Panel (a) depicts this for the case of sub-quadratic learning. As Theorem 1 (i) establishes, the risk-dominant boundary is selected so limit play is history independent, and does not depend on the shock’s realization. Panel (b) depicts the case of super-quadratic learning which rotates the boundary clockwise around the point (1/2,c1/2)(1/2,c-1/2). [^30] We have depicted this so that the state selects initial-play dominance i.e., $\lambda_{\infty}=1\iff\theta\geq c-\lambda_{0}$. We will later establish (Proposition 2 that initial-play dominance and risk-dominance span possible limit outcomes implementable via any learning speed $(\sigma_{t})_{t}$. More generally, for a given super-quadratic learning rate, this will lead to a (potentially non-linear) boundary between these two extremes. Hence, when learning is fast, the regime might now survive when the state is weak as long as the initial shock is small (region $A$); on the other hand, the regime might now also fail when the state is strong when the initial shock is large (region $B$).
@@
Figure 4 illustrates the forces underlying Theorem 1 (i) for the case where $\lambda_{0}<1/2$. Each panel corresponds to a different time. In each panel, the dotted line represents the state $\theta$, and the red curve is the distribution of posterior means $\mu_{t}$. The blue line is the threshold belief $\mu_{t}^{*}$, and the shaded area is the mass of agents $\lambda_{t}$ who take the risky action because their beliefs exceed the threshold. As time passes, agents learn and $\mu_{t}$ contracts around $\theta$. In Figure 4, learning and hence the speed of this contraction is slow. Consequently, the evolution of $\mu_{t}^{*}$ is relatively fast. As $t\to\infty$, $\mu_{t}^{*}$ converges to c(1/2)c-(1/2) and $\mu_{t}$ converges to a point mass on $\theta$ due to limit learning. Since $\theta\geq c-(1/2)$, aggregate play converges to the risk dominant (risky) action.
@@
Figure 5 illustrates the forces underlying Theorem 1 (ii) for the case where $\lambda_{0}<1/2$. With super-quadratic learning, the speed of contraction of $\mu_{t}$ is fast. Consequently, the evolution of $\mu_{t}^{*}$ is relatively slow. As $t\to\infty$, $\mu_{t}^{*}$ converges to a limit threshold $\mu_{\infty}^{*}$ that is bounded away from c(1/2)c-(1/2), and since $\theta\in(c-(1/2),\mu_{\infty}^{*})$, aggregate play converges to the non-risk dominant (safe) action.
@@
In particular, for any threshold $\mu^{*}$ between $c-\lambda_{0}$ and c(1/2)c-(1/2), there exists some learning process $\Sigma\geq 0$ such that:
$\lambda_{\infty}(\theta|\Sigma,\lambda_{0})=\mathbbm{1}(\theta\geq\mu^{*})\quad\text{a.e.}$
@@ 5 Transition dynamics
We have thus far focused on limit play. We now turn to studying the connection between the speed of learning and the time path of aggregate play. We will consider the iid signal case such that the learning process $\Sigma(\sigma)$ is parametrized by the standard deviation of each period’s signal $\sigma>0$. We focus on the iid case because it is a canonical model of learning, and makes results easier to state since $\Sigma$ is now parametrized by a scalar; nonetheless, we expect that the qualitative features of our results extend beyond this case. Within this iid learning environment, recall that Corollary 2 (i) implies that the limit action is risk dominant for all fundamentals and initial conditions. The following result shows that the path of aggregate play exhibits qualitatively different behavior depending on learning rates.
@@
Fix any $\lambda_{0}\in(0,1)$ and $\theta$ between c1/2c-1/2 and $c-\lambda_{0}$. Define $\overline{\beta}:=1+|c-(1/2)-\theta|/2$. For any $\epsilon>0$, $\alpha\in(0,1)$, and $\beta\in(1,\overline{\beta})$, there exists $\underline{\sigma},\overline{\sigma}>0$ such that:
@@
(b) $\sigma=0.01$: sudden phase transition
Proposition 3 has a subtle connection to the extant literature on coordination games. A common theme in this literature is that heterogeneity—in preferences, beliefs, etc.—across players can induce uniqueness in static supermodular games (Morris and Shin, 2006). A dynamic version of this insight is that with lots of hetrogeneity, as aggregate incentives shift, the set of equilibria vary smoothly—this is analogous to Part (i) of Proposition 3; by contrast, with little heterogeneity, the set of equilbiria can vary discontinuously—this is analogous to Part (ii) of Proposition 3. The key insight of Proposition 3 is that learning rates induce time-varying hetrogeneity in beliefs which, in turn, matters for transition dynamics. [^33] Koh and Morris (2022) relates heterogeneity to the speed of contagion in a networked population. This can be equivalently interpreted (see, e.g., Morris (1997)) as best-response dynamics or interim deletion of strictly dominated strategies where each player’s beliefs and higher-order beliefs are encoded by the network. However, Koh and Morris (2022) study speed of contagion for graphs/beliefs which are held fixed i.e., in the incomplete information game interpretation, there is no learning.
@@ 6 Concluding Remarks
We have developed a simple but novel model of inertial coordination games. This allowed us to combine and unify insights from the literature on best-response dynamics in complete-information coordination games (Crawford, 1995) and global games (Carlsson and Van Damme, 1993). We think that inertial coordination games serve as a natural description of many ‘real world’ coordination environments with both shrinking uncertainty about fundamentals (e.g., solvency of a bank, strength of a currency regime, quality of a networked product) as well as an endogenously changing state shaped by past play (e.g., past withdrawals, short positions, past sales). At the same time, we also showed that inertial coordination games share much of the game-theoretic logic underpinning coordination and learning in games. [^34] We show in Appendix II that for a fixed speed of learning (as a function of physical time), limit equilibrium predictions coincide as inertia vanishes (gap between time periods shrink). In this regard, inertial coordination games generalizes the setting with contemporaneous coordination motives. Nonetheless for any fixed level of inertia, the sub/super-quadratic boundary continues to hold. In this regard, we offer a different perspective on ‘history versus expectations’ (Krugman, 1991) in economic modeling—in the presence of inertia, expectations about history shape coordination outcomes. In particular, the speed at which players grow confident is a crucial determinant for whether coordination succeeds—the risk-dominant action is played in the limit if and only if posterior precisions grow sub-quadratically (Theorem 1). This delivered a suite of rich predictions of whether and to what extent initial shocks can propagate. In the super-quadratic regime, we showed that any profile of limit play between risk- and initial play-dominance can be rationalized by some learning speed. In the sub-quadratic regime, we further identified qualitatively different behavior in the path of aggregate play (Proposition 3): when learning is slow, transition dynamics are gradual; when learning is fast, sudden.
@@ A.1 System Dynamics
Conditional on $\theta$, the time-(t1)(t-1) distribution of posterior means is $\mu_{j,t-1}|\theta\sim N(\theta,\eta_{t-1}^{2})$. Since agents’ information sets are independent conditional on $\theta$, this implies that, under agent $i$’s time-$t$ posterior belief, the marginal distribution of time-(t1)(t-1) posterior means is $\mu_{j,t-1}|\mathcal{F}_{it}\sim N(\mu_{it},\eta_{t}^{2}+\eta_{t-1}^{2})$. Hence, by the inductive step, $i$ takes $a_{it}=1$ if and only if
@@
Let $F(\mu_{it})$ be the RHS function. Since $F$ is continuous and strictly increasing in $\mu_{it}$, and $F(c)>c$ and F(c1)<cF(c-1)<c hold, $F(\mu_{it})=c$ has a unique solution $\mu_{t}^{*}\in(c-1,c)$. Hence, there exists a unique cutoff $\mu_{t}^{*}$, defined implicitly by ($\mu^{*}$-LoM) at time $t$, such that $i$ takes $a_{it}=1$ if and only if $\mu_{it}\geq\mu_{t}^{*}$.
Finally, since $\mu_{it}|\theta\sim N(\theta,\eta_{t}^{2})$, we see that $\lambda_{t}(\theta)=\mathbb{P}_{\theta}(\mu_{it}\geq\mu_{t}^{*})=\Phi\left(\frac{\theta-\mu_{t}^{*}}{\eta_{t}}\right)$ and hence ($\lambda$-LoM) holds at time $t$, as desired.
@@
Assume $\mu_{t-1}^{*}=c-(1/2)$. Since $\mu_{t}^{*}:=c-(1/2)$ solves ($\mu^{*}$-LoM) and it has a unique solution, c(1/2)c-(1/2) is a steady state. To see that it’s the unique steady state, assume $\mu_{t-1}^{*}=\mu_{t}^{*}$. Then, ($\mu^{*}$-LoM) reduces to $\mu_{t}^{*}+\Phi(0)=c$ and hence $\mu_{t}^{*}=c-(1/2)$, as desired.
@@
$\mu_{t+1}^{*}>\mu_{t}^{*}\Leftrightarrow\quad\Phi\left(\frac{\mu_{t+1}^{*}-\mu_{t}^{*}}{\sqrt{\eta_{t}^{2}+\eta_{t+1}^{2}}}\right)>\frac{1}{2}\Leftrightarrow\quad\mu_{t+1}^{*}<\mu_{ss}^{*}.$ (From ($\mu^{*}$-LoM))
@@
    $|\mu^{*}_{t+1}-\mu_{t}^{*}|$ is decreasing in tt. [^35] Note that parts (i),(ii), and (iii) of this Lemma hold without Assumption 1.
@@
Part (iii). Note that ($\mu^{*}$-LoM) implies $|\mu^{*}_{t+1}-\mu_{t}^{*}|=\sqrt{\eta_{t}^{2}+\eta_{t+1}^{2}}\left|\Phi^{-1}(c-\mu_{t+1}^{*})\right|.$ Note that $\sqrt{\eta_{t}^{2}+\eta_{t+1}^{2}}$ is non-increasing in $t$. First, assume $\lambda_{0}>\frac{1}{2}$. Then, $(c-\mu_{t+1}^{*})>\frac{1}{2}$ decreasing in $t$ implies $\Phi^{-1}(c-\mu_{t+1}^{*})>0$ is decreasing in $t$, and therefore $|\Phi^{-1}(c-\mu_{t+1}^{*})|$ is decreasing in $t$. Hence, $|\mu^{*}_{t+1}-\mu_{t}^{*}|$ is also decreasing in $t$.
@@
Part (iv). Wlog, assume $\lambda_{0}>\frac{1}{2}$. First, we show $(\theta-\mu_{t}^{*})/\eta_{t}\to$-$\infty$ when $\theta<\mu_{\infty}^{*}$. Since $((\theta-\mu_{t}^{*})/\eta_{t})_{t}$ is decreasing and becomes negative for sufficiently large $t$, it should either converge to some negative constant or diverge to -$\infty$. Since $\lim_{t\to\infty}(\theta-\mu_{t}^{*})<0$, $\lim_{t\to\infty}(\theta-\mu_{t}^{*})/\eta_{t}=$-$\infty$ must hold by limit learning.
@@
Part (i). Subtracting ($\mu^{*}$-LoM) at time t1t-1 from ($\mu^{*}$-LoM) at time $t$ yields:
@@
Wlog, assume $\lambda_{0}>(1/2)$, then $\gamma_{1}>0$. Suppose for a contradiction that $\gamma_{2}<0$: then $\gamma_{1}<\gamma_{2}<0$, which is a contradiction since $\gamma_{1}>0$ by definition. Hence, $\gamma_{1}\geq\gamma_{2}\geq 0$. By induction, $(\gamma_{t})_{t}$ is a monotone decreasing sequence bounded below by 0, so it converges: $\gamma_{t}\downarrow\gamma_{\infty}\geq 0$, as desired. ∎
@@
Part (i). Fix any $x>0$. For any $A^{\prime}>A>0$, note that $A^{\prime}F_{A^{\prime}}(x)-AF_{A}(x)=\Phi(F_{A}(x))-\Phi(F_{A^{\prime}}(x))$. Suppose for a contradiction that $F_{A^{\prime}}(x)\geq F_{A}(x)$. Since $\Phi$ is increasing, the above implies $A^{\prime}F_{A^{\prime}}(x)\leq AF_{A}(x)$, which is a contradiction since $A^{\prime}>A$.
@@
$Ay^{\prime}=\varphi(x)-\varphi(y)y^{\prime}\implies y^{\prime}=\frac{\varphi(x)}{A+\varphi(y)}>0,$
@@
Then $A_{\gamma_{1}}(x)\uparrow\bar{A}^{*}$ follows from continuity of $A_{\gamma_{1}}(\cdot)$. ∎